# HSIGene: A Foundation Model For Hyperspectral Image Generation

### v1

@@ Abstract
Hyperspectral image (HSI) plays a vital role in various fields such as agriculture and environmental monitoring. However, due to the expensive acquisition cost, the number of hyperspectral images is limited, degenerating the performance of downstream tasks. Although some recent studies have attempted to employ diffusion models to synthesize HSIs, they still struggle with the scarcity of HSIs, affecting the reliability and diversity of the generated images. Some studies propose to incorporate multi-modal data to enhance spatial diversity, but the spectral fidelity cannot be ensured. In addition, existing HSI synthesis models are typically uncontrollable or only support single-condition control, limiting their ability to generate accurate and reliable HSIs. To alleviate these issues, we propose HSIGene, a novel HSI generation foundation model which is based on latent diffusion and supports multi-condition control, allowing for more precise and reliable HSI generation. To enhance the spatial diversity of the training data while preserving spectral fidelity, we propose a new data augmentation method based on spatial super-resolution, in which HSIs are upscaled first, and thus abundant training patches could be obtained by cropping the high-resolution HSIs. In addition, to improve the perceptual quality of the augmented data, we introduce a novel two-stage HSI super-resolution framework, which first applies RGB bands super-resolution and then utilizes our proposed Rectangular Guided Attention Network (RGAN) for guided HSI super-resolution. Experiments demonstrate that the proposed model is capable of generating a vast quantity of realistic HSIs for downstream tasks such as denoising and super-resolution. The code and models are available at https://github.com/LiPang/HSIGene.
@@ I Introduction
Hyperspectral image (HSI) is captured across a continuous range of wavelengths, providing detailed information on the spectral characteristics of different materials. HSI plays a crucial role in various applications such as remote sensing [1, 2, 3], medical [4, 5, 6] and agriculture [7, 8, 9]. In recent years, with the rapid advancement of artificial intelligence, deep learning (DL) techniques have been widely adopted across various HSI applications, including classification [10, 11, 12, 13], denoising [14, 15, 16, 17], super-resolution [18, 19] and so on. However, owing to the high cost associated with hyperspectral data acquisition, the number of high-quality HSIs is limited, posing significant challenges to the application of DL in HSI processing tasks.
@@
To alleviate the limitation of the scarcity of HSIs, a promising alternative approach involves the synthetic generation of HSIs using deep generative models. Deep generative models, such as Variational Autoencoders (VAEs) [20, 21], Generative Adversarial Networks (GANs) [22, 23], and Diffusion Models (DMs) [24, 25], have shown remarkable success in generating realistic data across various domains. Among them, diffusion models have gained increasing popularity, since diffusion models have been shown to achieve excellent generation results and can avoid some of the common drawbacks of VAEs and GANs, such as mode collapse in GANs and posterior collapse in VAEs. Furthermore, the technique of Latent Diffusion Models (LDMs) [26] combines the power of diffusion models with the efficiency of latent space representations, significantly reducing computational burden while maintaining desirable visual fidelity. Owing to the impressive synthesis capability, the application of generative models to hyperspectral imaging is an area of growing interest.
@@
Figure 1: The comparison of the existing HSI augmentation methods. UBF [27] alleviates the issue of the data sacrifice by performing spectral super-resolution on external RGB images while our method generates new HSI samples by performing spatial super-resolution on existing real HSIs.
@@
Figure 2: Visualization results of our proposed HSIGene in different situations. (a) Unconditional generation results. (b) Generation results under single condition. (c) Generation results under multiple conditions.
@@
Recently, some studies have attempted to synthesize HSIs with diffusion models [27, 28]. One representative work, Unmixing before Fusion (UBF) [27], synthesizes hyperspectral data in the abundance domain and incorporates RGB images to increase the diversity of training HSIs. Specifically, an unmixing network is first trained on real HSIs to estimate endmembers and abundances. After that, as shown in Fig. 1(a), the abundances of external RGB images are inferred utilizing the unmixing network. Finally, a fusion-based generative model is trained to synthesize abundance maps, and new HSI samples are generated by combining the synthesized abundance maps with the estimated endmembers. Although the model can generate HSIs, due to the inherent spectral and content differences between hyperspectral and RGB data, the fidelity of the spectral profiles of the generated HSIs cannot be guaranteed. Instead, UnmixDiff [28], directly trains the generative model with the abundance maps of real HSIs. However, UnmixDiff still suffers from the scarcity of high-quality HSIs and limited model size, limiting the diversity of the generated HSIs. In addition, UBF and Unmixdiff both synthesize abundance maps instead of HSIs directly, resulting that the quality of the generated spectra is highly limited by the performance of unmixing network. Moreover, only a single condition can be incorporated to control the synthesis process, and thus Unmixdiff cannot guarantee the accuracy of generated HSIs.
@@
UBF Synthetic 0
UnmixDiff Real 2
@@
To alleviate these issues, we propose a novel HSI synthesis foundation model, HSIGene, which is based on latent diffusion models and supports multi-condition controllable generation. As illustrated in Fig. 2, our model is capable of converting multiple conditions into synthesized HSIs. To alleviate the scarcity of high-quality HSIs and enhance the spatial diversity of training samples, we perform spatial super-resolution (SR) on existing hyperspectral data as shown in Fig. 1(b). Abundant HSI samples could be obtained by cropping the upscaled HSIs instead of the original real data. Since there are rich land covers inside a real HSI, the cropping patches of the upscaled HSIs exhibit various content, leading to better generation performance and generalization capabilities. Considering the challenges associated with real-world hyperspectral super-resolution (i.e. the lack of ground-truth high-resolution HSIs for training networks), we adopt a two-stage super-resolution framework that leverages the power of abundant high-resolution RGB images and diffusion models to improve the perceptual quality of the augmented data. Specifically, we first collect a large dataset of high-resolution RGB images that closely resemble real hyperspectral data. Utilizing the collected high-resolution RGB images, a diffusion model-based RGB super-resolution network (DSRNet) is trained and is then used to super-resolve the RGB bands of real HSIs, resulting in high-resolution RGB bands. Finally, with high-resolution RGB bands as guidance, a novel rectangular guided attention network (RGAN) is proposed to obtain high-resolution HSIs. A more detailed comparison of our model and existing works is demonstrated in Tab. I.
Overall, our contributions could be summarized as follows.
  • We propose a foundation model namely HSIGene for hyperspectral image generation. Our model can generate high-quality HSIs under various control conditions such as sketch and segmentation. As far as we know, HSIGene is the first and largest generative model which supports multiple control conditions for HSI generation.
@@
  • To improve the perceptual quality of augmented data, we propose a two-stage framework for HSI super-resolution including RGB bands super-resolution and RGB-guided HSI super-resolution. In addition, a novel rectangular guided attention network (RGAN) is proposed to fully transfer the detailed supplementary information from RGB modality to HSI modality.
@@
The rest of the paper is organized as follows: In Section II, we review the related work, focusing on the advancements in generative models, HSI synthesis and HSI super-resolution. Section III presents our proposed HSIGene in detail, outlining the synthesis framework and the architecture of guided super-resolution network. Section IV provides a comprehensive analysis of the synthesis results and Section V summarizes our contributions.
II Related Works
II-A Deep Generative Model
Deep generative models have emerged as powerful tools for generating samples from complex data distributions, with significant applications across various domains. These models could be primarily categorized into three types: Variational Autoencoders (VAEs) [20, 21], Generative Adversarial Networks (GANs) [22, 23], and Diffusion Models (DMs) [24, 25].
VAEs [20, 21] learn to encode data into a latent space by maximizing the evidence lower bound (ELBO). They consist of an encoder-decoder structure, where the encoder maps input data to a latent space, and the decoder reconstructs the data from the latent representation. VAEs are known for their ability to generate smooth variations of data by sampling from the latent space, but they often suffer from blurry outputs due to the inherent trade-off between reconstruction fidelity and the smoothness of the latent space [29]. GANs [22, 23] are typically composed of a generator and a discriminator, and the two networks are trained in a competitive way to continually improve each other with the aim of achieving a Nash equilibrium. GANs have achieved remarkable success in generating high-quality realistic images, but they are often difficult to train due to issues such as mode collapse and instability during training [30, 31].
@@ II-A Deep Generative Model
Recently, diffusion models [24, 25] have become a popular choice for data synthesis in various domains. Diffusion models gradually transform simple noise distributions into complex data distributions through a series of iterative denoising steps. Among these methods, Latent Diffusion Models (LDMs) [26, 32] have gained attention for their ability to operate in a compressed latent space, significantly reducing computational costs while maintaining high generation quality. In addition, the cross-attention layers in LDMs enable the models to generate content by conditioning on inputs such as text or bounding boxes. Over the past two years, various Controllable Diffusion Models (CDMs) have made tremendous progress, enabling the generation of highly realistic images based on various inputs, such as text, sketches, or specific attributes. For example, T2I-Adapter [33] proposes a lightweight and efficient model to enhance the controllability of pre-trained text-to-image diffusion models without altering original network structure. The T2I-Adapters are trained to align external control signals with the internal knowledge of the T2I models, enabling more granular control over attributes such as color and structure in the generated images. Uni-ControlNet [34] integrates diverse local and global controls into text-to-image diffusion models through the use of two additional adapters, offering flexible and composable control over image generation. There are also some generative works in the remote sensing field. DiffusionSat [35] employs a novel 3D ControlNet to enable more flexible and high-quality generation of satellite images for various applications. CRS-Diff [36] integrates the capabilities of diffusion models with advanced control mechanisms, supporting multiple control inputs, including text, metadata, and image conditions, to guide the generation process. However, HSI synthesis with multiple conditions remains unexplored.
II-B Hyperspectral Image Synthesis
@@
Owing to limited high-quality HSI data and the characteristics of high dimensionality, only a few methods that focus on HSI synthesis with deep generative models [27, 28] are proposed. Unmixing before Fusion (UBF) [27] introduces a novel paradigm for synthesizing HSIs by leveraging unmixing across multi-source data followed by fusion-based synthesis. In the method, an unmixing network is first trained on HSI data to extract endmembers and abundance maps. Using the trained unmixing model, the abundance from unpaired RGB data is inferred to train diffusion models. Finally, by fusing the estimated endmembers with the synthetic abundances generated by diffusion models, the approach effectively addresses the data scarcity issue in HSI researches. Due to the inherent spectral and content differences between hyperspectral and RGB data, the model, while capable of generating HSIs, cannot ensure the fidelity of the spectral profiles of the generated images. UnmixDiff [28], similar to UBF, also generates HSIs by synthesizing abundance maps, which are then fused with estimated spectral endmembers to synthesize HSIs. While trained directly on real HSIs, UnmixDiff still suffers the scarcity of high-quality HSIs and only supports generation with single condition.
@@
II-C Hyperspectral Image Super-resolution
The majority of existing HSI-SR approaches are based on Convolutional Neural Networks (CNN) and Transformers. For example, GDRRN [37] employs a grouped recursive module within a deep neural network to enhance the spatial resolution of HSIs while minimizing spectral distortion. MCNet [38] utilizes a novel mixed convolutional module (MCM) that combines 2D and 3D convolutions to effectively extract spatial and spectral features for HSI super-resolution. Bi-3DQRNN [39] integrates a 3D convolutional module with a bidirectional quasi-recurrent pooling module to capture spatial-spectral structures and global spectral correlations. SSPSR [40] employs a group convolution strategy with shared parameters and a progressive upsampling framework for super-resolution of hyperspectral imagery. ESSAformer [41] incorporates self-attention mechanism to further capture spatial-spectral information and long-range dependencies. However, these methods ignore the ill-posed nature of super-resolution task and tend to generate blurry outputs. Recently, a few approaches propose to employ diffusion models to enhance the super-resolution performance. DMGASR [42] integrates a Group-Autoencoder and a diffusion model to enhance the spatial resolution of HSIs while preserving spectral correlations. S2CycleDiff [43] leverages a conditional cycle-diffusion process and spatial/spectral guided pyramid denoising to enhance both spatial details and spectral accuracy. While these models demonstrate promising performance when applied to datasets similar to training data, such approaches suffer the scarcity of HSI data and could exhibit poor generalization in real-world settings.
III Methods
III-A Dataset Description
III-A1 HSI data collection
In the process of compiling the training dataset, we encountered a variety of data sources with different qualities and characteristics. However, not all data are suitable for inclusion in our study. Specifically, we choose to exclude certain datasets due to their low resolution (e.g., Hyspecnet-11k [44], WHU-OHS [45]), high noise levels (e.g., Hyperion data [^1] https://www.usgs.gov/centers/eros/science/usgs-eros-archive-earth-observing-one-eo-1-hyperion), small size (e.g., Salinas [^2] https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes), and the obscurity of semantic information. Finally, five HSI datasets, including Xiongan [46], Chikusei [47], DFC2013333https://hyperspectral.ee.uh.edu/?page_id=459, DFC2018444https://hyperspectral.ee.uh.edu/?page_id=1075 and Heihe [48, 49], are employed as the training set of the generative models. A more detailed description of these datasets is provided in Table II. However, the scarcity of high-quality HSIs might hinder the ability of our model to generate images from diverse conditions. Therefore, we propose a novel data augmentation strategy by performing spatial super-resolution to improve the generalization performance of our model, and more details are illustrated in Sec. III-C.

### v2

@@ Abstract
Hyperspectral image (HSI) plays a vital role in various fields such as agriculture and environmental monitoring. However, due to the expensive acquisition cost, the number of hyperspectral images is limited, degenerating the performance of downstream tasks. Although some recent studies have attempted to employ diffusion models to synthesize HSIs, they still struggle with the scarcity of HSIs, affecting the reliability and diversity of the generated images. Some studies propose to incorporate multi-modal data to enhance spatial diversity, but spectral fidelity cannot be ensured. In addition, existing HSI synthesis models are typically uncontrollable or only support single-condition control, limiting their ability to generate accurate and reliable HSIs. To alleviate these issues, we propose HSIGene, a novel HSI generation foundation model which is based on latent diffusion and supports multi-condition control, allowing for more precise and reliable HSI generation. To enhance the spatial diversity of the training data while preserving spectral fidelity, we propose a new data augmentation method based on spatial super-resolution, in which HSIs are upscaled first, and thus abundant training patches could be obtained by cropping the high-resolution HSIs. In addition, to improve the perceptual quality of the augmented data, we introduce a novel two-stage HSI super-resolution framework, which first applies RGB bands super-resolution and then utilizes our proposed Rectangular Guided Attention Network (RGAN) for guided HSI super-resolution. Experiments demonstrate that the proposed model is capable of generating a vast quantity of realistic HSIs for downstream tasks such as denoising and super-resolution. The code and models are available at https://github.com/LiPang/HSIGene.
@@ 1 Introduction
Hyperspectral image (HSI) is captured across a continuous range of wavelengths, providing detailed information on the spectral characteristics of different materials. HSI plays a crucial role in various applications such as remote sensing [1, 2], medical [3, 4] and agriculture [5, 6]. In recent years, with the rapid advancement of artificial intelligence, deep learning (DL) techniques have been widely adopted across various HSI applications, including classification [7, 8], denoising [9, 10, 11], super-resolution [12, 13] and so on. However, owing to the high cost associated with hyperspectral data acquisition, the number of high-quality HSIs is limited, posing significant challenges to the application of DL in HSI processing tasks.
@@
To alleviate the limitation of the scarcity of HSIs, a promising alternative approach involves the synthetic generation of HSIs using deep generative models. Deep generative models, such as Variational Autoencoders (VAEs) [14], Generative Adversarial Networks (GANs) [15], and Diffusion Models (DMs) [16, 17], have shown remarkable success in generating realistic data across various domains. Among them, diffusion models have gained increasing popularity, since they have achieved excellent generation results and can avoid some of the common drawbacks of VAEs and GANs, such as mode collapse in GANs and posterior collapse in VAEs. Furthermore, the technique of Latent Diffusion Models (LDMs) [18] combines the power of diffusion models with the efficiency of latent space representations, significantly reducing computational burden while maintaining desirable visual fidelity. Owing to its impressive synthesis capability, applying generative models to hyperspectral imaging is an area of growing interest.
@@
Figure 1: A schematic comparison of the existing HSI augmentation methods. UBF [19] alleviates the issue of the data sacrifice by performing spectral super-resolution on external RGB images, resulting in data with similar content to RGB images but with less authentic spectral profiles. In contrast, our method performs spatial super-resolution on existing real HSIs, ensuring that both the content and spectral distribution are consistent with real HSIs.
@@
Figure 2: Visualization results of our proposed HSIGene in different situations. When provided with more conditions, they can complement each other to achieve more accurate generation. (a) Unconditional generation results. (b) Generation results under single condition and multiple conditions.
@@
Recently, some studies have attempted to synthesize HSIs with diffusion models [19, 20]. One representative work, Unmixing before Fusion (UBF) [19], synthesizes hyperspectral data in the abundance domain and incorporates RGB images to increase the diversity of training HSIs. Specifically, an unmixing network is first trained on real HSIs to estimate endmembers and abundances. After that, as shown in Fig. 1(a), the abundances of external RGB images are inferred utilizing the unmixing network. Finally, a fusion-based generative model is trained to synthesize abundance maps, and new HSI samples are generated by combining the synthesized abundance maps with the estimated endmembers. Although the model can generate HSIs, due to the inherent spectral and content differences between hyperspectral and RGB data, the fidelity of the spectral profiles of the generated HSIs cannot be guaranteed. Instead, UnmixDiff [20], directly trains the generative model with the abundance maps of real HSIs. However, UnmixDiff still suffers from the scarcity of high-quality HSIs and limited model size, limiting the diversity of the generated HSIs. In addition, UBF and Unmixdiff both synthesize abundance maps instead of HSIs directly, resulting in the quality of the generated spectra being highly limited by the performance of the unmixing network. Moreover, only a single condition can be incorporated to control the synthesis process, and thus Unmixdiff cannot guarantee the accuracy and reliability of the generated HSIs.
@@
UBF [19] Synthetic 0
UnmixDiff [20] Real 2
@@
To alleviate these issues, we propose a novel HSI synthesis foundation model, HSIGene, which is based on latent diffusion models and supports multi-condition controllable generation. As illustrated in Fig. 2, our model is capable of generating HSIs both unconditionally and under the guidance of one or more conditions. Multiple conditions can be combined to provide more comprehensive information, resulting in more precise generation. To alleviate the scarcity of high-quality HSIs and enhance the spatial diversity of training samples, we perform spatial super-resolution (SR) on existing hyperspectral data as shown in Fig. 1(b). Abundant HSI samples could be obtained by cropping the upscaled HSIs instead of the original real data. Since there are rich land covers inside a real HSI, the cropping patches of the upscaled HSIs exhibit various content, leading to better generation performance and generalization capabilities. Considering the challenges associated with real-world hyperspectral super-resolution (i.e. the lack of ground-truth high-resolution HSIs for training networks), we adopt a two-stage super-resolution framework that leverages the power of abundant high-resolution RGB images and diffusion models to improve the perceptual quality of the augmented data. Specifically, we first collect a large dataset of high-resolution RGB images that closely resemble real hyperspectral data by utilizing the latitude and longitude information. Utilizing the collected high-resolution RGB images, a diffusion model-based RGB super-resolution network (DSRNet) is trained and is then used to super-resolve the RGB bands of real HSIs, resulting in high-resolution RGB bands. Finally, with high-resolution RGB bands as guidance, a novel rectangular guided attention network (RGAN) is proposed to obtain high-resolution HSIs. A more detailed comparison of our model and existing works is demonstrated in Tab. I.
Overall, our contributions can be summarized as follows.
  • We propose a foundation model namely HSIGene for hyperspectral image generation. Our model can generate high-quality HSIs under various control conditions such as sketch and segmentation. As far as we know, HSIGene is the first and largest HSI generative model, supporting multiple control conditions.
@@
  • To improve the perceptual quality of augmented data, we propose a two-stage framework for HSI super-resolution including RGB bands super-resolution and RGB-guided HSI super-resolution. Additionally, a novel rectangular guided attention network (RGAN) is proposed to fully transfer the detailed supplementary information from the RGB modality to the HSI modality.
@@
The rest of the paper is organized as follows: In Section 2, we review the related work on the generative models, HSI synthesis and HSI super-resolution. Section 3 presents our proposed HSIGene in detail, outlining the synthesis framework and the architecture of a guided super-resolution network. Section 4 provides comprehensive experimental results and Section 5 summarizes our work.
2 Related Works
2.1 Diffusion Model
@@
Recently, diffusion models [16, 17] have emerged as a powerful tool for generating samples from complex data distributions, with significant applications across various domains. Diffusion models gradually transform simple noise distributions into complex data distributions through a series of iterative denoising steps. Among these methods, Latent Diffusion Models (LDMs) [18] have gained attention for their ability to operate in a compressed latent space, significantly reducing computational costs while maintaining high generation quality. Besides, the cross-attention layers in LDMs enable the models to generate content by conditioning on inputs such as text or bounding boxes. Over the past two years, various Controllable Diffusion Models (CDMs) have been proposed, enabling the generation of highly realistic images based on various inputs, such as text, sketches, or specific attributes. For example, T2I-Adapter [21] proposes a lightweight and efficient model to enhance the controllability of pre-trained text-to-image diffusion models without altering the original network structure. Uni-ControlNet [22] integrates diverse local and global controls into text-to-image diffusion models through the use of two additional adapters, offering flexible and composable control over image generation. There are also some generative works in the remote sensing field. DiffusionSat [23] employs a novel 3D ControlNet to enable a more flexible and high-quality generation of satellite images for various applications. CRS-Diff [24] integrates the capabilities of diffusion models with advanced control mechanisms, supporting multiple control inputs, including text, metadata, and image conditions, to guide the generation process. However, HSI synthesis with multiple conditions remains unexplored.
2.2 Hyperspectral Image Synthesis
@@
Owing to limited high-quality HSI data and the characteristics of high dimensionality, only a few methods that focus on HSI synthesis with deep generative models [19, 20] are proposed. Unmixing before Fusion (UBF) [19] introduces a novel paradigm for synthesizing HSIs by leveraging unmixing across multi-source data followed by fusion-based synthesis. In the method, an unmixing network is first trained on HSI data to extract endmembers and abundance maps. Using the trained unmixing model, the abundance from unpaired RGB data is inferred to train diffusion models. Finally, by fusing the estimated endmembers with the synthetic abundances generated by diffusion models, the approach effectively addresses the data scarcity issue in HSI research. Due to the inherent spectral and content differences between hyperspectral and RGB data, the model, while capable of generating HSIs, cannot ensure the fidelity of the spectral profiles of the generated images. UnmixDiff [20], similar to UBF, also generates HSIs by synthesizing abundance maps, which are then fused with estimated spectral endmembers to synthesize HSIs. While trained directly on real HSIs, UnmixDiff still suffers from the scarcity of high-quality HSIs and only supports generation with a single condition.
2.3 Hyperspectral Image Super-resolution
The majority of existing HSI-SR approaches are based on Convolutional Neural Networks (CNN) and Transformers. For example, GDRRN [25] employs a grouped recursive module within a deep neural network to enhance the spatial resolution of HSIs while minimizing spectral distortion. MCNet [26] utilizes a novel mixed convolutional module (MCM) that combines 2D and 3D convolutions to effectively extract spatial and spectral features for HSI super-resolution. Bi-3DQRNN [27] integrates a 3D convolutional module with a bidirectional quasi-recurrent pooling module to capture spatial-spectral structures and global spectral correlations. SSPSR [28] employs a group convolution strategy with shared parameters and a progressive upsampling framework for super-resolution of hyperspectral imagery. ESSAformer [29] incorporates a self-attention mechanism to further capture spatial-spectral information and long-range dependencies. However, these methods ignore the ill-posed nature of super-resolution tasks and tend to generate blurry outputs. Recently, a few approaches proposed to employ diffusion models to enhance the super-resolution performance. DMGASR [30] integrates a Group-Autoencoder and a diffusion model to enhance the spatial resolution of HSIs while preserving spectral correlations. S2CycleDiff [31] leverages a conditional cycle-diffusion process and spatial/spectral guided pyramid denoising to enhance both spatial details and spectral accuracy. While these models demonstrate promising performance when applied to datasets similar to training data, such approaches suffer the scarcity of HSI data and could exhibit poor generalization in real-world settings.
@@ 2.3 Hyperspectral Image Super-resolution
3 Methods
3.1 Dataset Description
3.1.1 HSI data collection
In the process of compiling the training dataset, we encountered a variety of data sources with different qualities and characteristics. However, not all data are suitable for inclusion in our study. Specifically, we choose to exclude certain datasets due to their low resolution (e.g., Hyspecnet-11k [32], WHU-OHS [33]), high noise levels (e.g., Hyperion data [^1] https://www.usgs.gov/centers/eros/science/usgs-eros-archive-earth-observing-one-eo-1-hyperion), small size (e.g., Salinas [^2] https://www.ehu.eus/ccwintco/index.php/Hyperspectral_Remote_Sensing_Scenes), and the obscurity of semantic information. Finally, five HSI datasets, including Xiongan [34], Chikusei [35], DFC2013333https://hyperspectral.ee.uh.edu/?page_id=459, DFC2018444https://hyperspectral.ee.uh.edu/?page_id=1075 and Heihe [36, 37], are employed as the training set of the generative models. A more detailed description of these datasets is provided in Table II. However, the scarcity of high-quality HSIs might hinder the ability of our model to generate images from diverse conditions. Therefore, we propose a novel data augmentation strategy by performing spatial super-resolution to improve the generalization performance of our model, and more details are illustrated in Sec. 3.3.